KyledaGreat:

Solve the following logarithmic equation. Express your answer as either an exact expression or a decimal approximation rounded to four decimal places. If there is no solution, indicate "No Solution (∅)." log_7(x−5) + log_7(x+3) = log_7(x−2) If you wish to enter log or ln, you must use the keypad. If there is more than one solution, separate your answers with a comma. Selecting a radio button will replace the entered answer value(s) with the radio button value. If the radio button is not selected, the entered answer is used. x = ____________________ No Solution (∅)

KyledaGreat:

@vocaloid

KyledaGreat:

is my answer right with \[x = \frac{ 3+\sqrt{61} }{ 2}\]

AZ:

(x-5)(x+3) = x-2 x^2 - 2x - 15 = x - 2 x^2 - 3x - 13 = 0 and the only positive solution is the one you listed so good job!